Genny is 14 years old. Her mother is 38. How many years ago was the mother exactly 5 times as old as Genny?
Answer by checkley77(12844)

5(14-x)=38-x
70-5x=38-x
-5x+x=38-70
-4x=-32
x=-32/-4
x=8 years ago is the ans.
Proof:
5(14-8)=38-8
5*6=30
30=30